IMPORTANCE OF RELATIVISTIC CORRECTIONS TO THE SPECTRUM OF WATER AND HYDROGEN SULPHIDE

\underline{P. BARLETTA}, J. TENNYSON, O.L. POLYANSKY, Department of Physics and Astronomy,University College London, Gower St., London WC1E 6BT, UK; and A.G. CSÁSZÁR , Department of Theoretical Chemistry, Eötvös University, H-1518 Budapest 112, PO Box 32, Hungary.

The importance of effects as small as minor relativistic corrections in theoretical predictions of energy levels can be as high as few wavenumbers [1]. Two-electron relativistic corrections to the ground-state electronic energy of water and hydrogen sulphide are determined at over 300 (250) geometries. The corrections include the two-electron Darwin term of the Coulomb–Pauli Hamiltonian, and the Gaunt and Breit corrections, calculated perturbationally using fourcomponent variational Dirac–Hartree–Fock wavefunctions. Fitted relativistic correction surfaces are constructed and used with an accurate ab initio nonrelativistic Born–Oppenheimer potential to calculate vibrational and rotational levels. The calculations suggest that these two-electron relativistic corrections, which have so far been neglected in rovibrational calculations on light molecules, have a substantial influence on the ro-vibrational levels [2]. The effects considered have markedly different characteristics for the stretching and bending levels.

[1] J. Tennyson, P. Barletta, M.A. Kosten, O.L. Polyansky and N.F. Zobov, Spectrochimica Acta A 58 (2002) 663.

[2] H.M. Quiney, P. Barletta, G. Tarczay, A.G. Császár, O.L. Polyansky and J.Tennyson, Chem. Phys. Lett. 344 (2001)